# SV4D: Dynamic 3D Content Generation with Multi-Frame and Multi-View Consistency

### v1

@@ 4 Experiments
Datasets. We evaluate SV4D-synthesized novel view videos and 4D optimization results on the synthetic datasets ObjaverseDy, Consistent4D [9], and the real-world videos dataset DAVIS [38; 39; 4]. Consistent4D dataset contains some dynamic objects collected from Objaverse [9; 8]. We excluded these objects from our training set to make a fair comparison. We used the same input video and evaluated views as Consistent4D. DAVIS dataset consists of single-view videos in the wild with corresponding foreground masks. For qualitative evaluation, we selected single-view videos with static cameras.
Metrics. We use the SV4D model to generate multiple novel view videos corresponding to the trajectories of the ground truth camera in the evaluation datasets. We compare each generated image with its corresponding ground-truth, in terms of Learned Perceptual Similarity (LPIPS [67]) and CLIP-score (CLIP-S) to evaluate visual quality. In addition, we evaluate the video coherence by reporting FVD [52], a video-level metric commonly used in video generation tasks. We calculate FVD with different ways (see Fig. 8):
  • FVD-F: calculate FVD over frames at each view.
  • FVD-V: calculate FVD over views at each frame.
  • FVD-Diag: calculate FVD over the diagonal images of the image matrix.
  • FV4D: calculate FVD over all images by scanning them in a bidirectional raster order.
Baselines. For novel view video synthesis, we compare SV4D with several recent methods capable of generating multiple novel view videos from a single-view video, including SV3D [54], Diffusion2 [61], STAG4D [61]. It is to be noted that all of these methods are inference-only techniques, and do not involve direct training in the 4D space like our method. We run SV3D to generate multi-view images at each video frame separately. Diffusion2 has not released their code yet. STAG4D used Zero123++ [44] as the multi-view generator, which fixed the view angle, and hence the novel views generated from STAG4D cannot be changed to be consistent with the views evaluated. We reproduced Diffusion2 and STAG4D with SV3D as the multi-view generator. SV3D has been shown to generate more consistent 3D results than Zero123++, so this serves as a better baseline. For 4D generation, we compare SV4D with other methods that can generate 4D representations, including Consistent4D [18], STAG4D [66], DreamGaussian4D (DG4D) [42], GaussianFlow [10], 4DGen [64], Efficient4D [37].
@@
SV3D [54] 0.129 0.925 989.53
Diffusion2 [61] 0.171 0.919 1430.35
STAG4D [66] 0.131 0.929 861.88
@@
Consistent4D [18] 0.160 0.87 1133.93
STAG4D [66] 0.126 0.91 992.21
Efficient4D [37] 0.140 0.92 -
4DGen [64] 0.130 0.89 -
DG4D [42] 0.160 0.87 -
GaussianFlow [10] 0.140 0.91 -
@@
Table 3: Evaluation of Novel View Video Synthesis on the ObjaverseDy Dataset. SV4D can get superior performance in video frame consistency and multi-view consistency. Our reproduced version.
@@
SV3D [54] 0.131 0.920 729.67 375.49 526.78 690.49
Diffusion2 [61] 0.152 0.905 1067.78 442.18 612.82 977.19
STAG4D [66] 0.133 0.917 652.43 469.07 636.83 546.56
@@
Consistent4D [18] 0.165 0.896 880.54 488.38 741.52 871.95
STAG4D [66] 0.158 0.860 929.10 453.62 663.50 1003.16
DreamGaussian4D [42] 0.152 0.897 697.80 450.58 615.68 638.15
@@ 4.1 Quantitative Comparison
Figure 6: Visual Comparison of Generated 4D Outputs. By leveraging the consistent multi-view videos generated by SV4D, we can learn higher-quality 4D assets compared to the prior works with SDS-based losses. Our results are more detailed, consistent, and faithful to the input videos.
@@ 4.2 Visual Comparison
4D Generation. We compare our generated 4D results with the prior methods in Fig. 6. For each video, we render the 4D outputs at two different timestamps and two novel views. Consistent4D and STAG4D often produce blurry outputs with inconsistent geometry and texture. DreamGaussian4D can generate finer texture details but still suffers from flickering artifacts and sometimes creates inaccurate geometry. Moreover, all these methods rely on the computationally expensive SDS loss, which is prone to spatially incoherent results and over-saturated texture. On the contrary, SV4D optimizes 4D assets using purely photometric and geometric losses, resulting in smoother videos with realistic and faithful object appearance.
@@ 4.3 User Study
SV3D [54] 700.72 831.18 656.59 780.99
Diffusion2 [61] 896.98 890.70 801.99 1093.24
STAG4D [66] 765.79 669.28 601.23 659.79
@@ 4.4 Ablative Analyses
SV4D Sampling Is Better Than Off-the-shelf Interpolation Method. Extending anchor frames to the full $V{\mkern-2.0mu\times\mkern-2.0mu}F$ image matrix is not trivial. In Table 6, we compare our SV4D sampling (ours) and the off-the-shelf interpolation method AMT [25]. We also show the results of anchor frames as a reference (SV4D (anchor)). The model with SV4D sampling outperforms AMT across all metrics. Fig. 8 validates this observation, where we can see that the SV4D sampling can synthesize distinct images while the results of the off-the-shelf interpolation method have a lot of motion blur. Table 6 also shows two intuitive sampling strategies, independent and interleaved sampling, which generate the videos with worse consistency. See the Appendix for details on these two sampling strategies. These results validate that the SV4D sampling is effective.
@@ 5 Conclusion
Acknowledgments. We thank Prem Akkaraju, Christian Laforte, Adam Letts, Josephine Parquet, Savannah Martin, Cathy Lee, Amir Etemadieh, Patricia Wood, Nadiia Klokova, Shan Shan Wong, Ryan Holeman, Jerry Chi, Javad Azimi for their help in various aspects of the model development and release. We also thank Eric Courtemanche for his help with visual results.

### v2

@@ 4 Experiments
Datasets. We evaluate SV4D-synthesized novel view videos and 4D optimization results on the synthetic datasets ObjaverseDy and Consistent4D (Deitke et al., 2023b). Consistent4D dataset contains dynamic objects collected from Objaverse (Deitke et al., 2023b; a). We excluded these objects from our training set to make a fair comparison. We used the same input video and evaluated views as Consistent4D. For the visual comparison, we also used single-view videos from the real-world videos dataset DAVIS (Perazzi et al., 2016; Pont-Tuset et al., 2017; Caelles et al., 2019).
Metrics. We use the SV4D model to generate multiple novel view videos corresponding to the trajectories of the ground truth camera in the evaluation datasets. We compare each generated image with its corresponding ground-truth, in terms of Learned Perceptual Similarity (LPIPS (Zhang et al., 2018b)) and CLIP-score (CLIP-S) to evaluate visual quality. In addition, we evaluate the video coherence by reporting FVD (Unterthiner et al., 2018), a video-level metric commonly used in video generation tasks. We calculate FVD with different ways (see Fig. 8): FVD-F: calculate FVD over frames at each view. FVD-V: calculate FVD over views at each frame. FVD-Diag: calculate FVD over the diagonal images of the image matrix. FV4D: calculate FVD over all images by scanning them in a bidirectional raster order.
Baselines. For novel view video synthesis, we compare SV4D with several recent methods capable of generating multiple novel view videos from a single-view video, including SV3D (Voleti et al., 2024), Diffusion2 (Yang et al., 2025), STAG4D (Zeng et al., 2024). It is to be noted that all of these methods are inference-only techniques, and do not involve direct training in the 4D space like our method. We run SV3D to generate multi-view images for each video frame separately. STAG4D used Zero123++ (Shi et al., 2023) as the multi-view generator, which fixed the view angle, and hence the novel views generated from STAG4D cannot be changed to be consistent with the views evaluated. We reproduced STAG4D with SV3D as the multi-view generator. SV3D has been shown to generate more consistent 3D results than Zero123++, so this serves as a stronger baseline. For reference, we also compare SV4D with the concurrent work 4Diffusion (Zhang et al., 2024). We do not report the FVD-F metric for 4Diffusion, as it only generates 8 frames, which is inconsistent with the number of frames used in our evaluation. Note that the FVD-F metric is highly sensitive to the number of frames. For 4D generation, we compare SV4D with other methods that can generate 4D representations, including Consistent4D (Jiang et al., 2024c), STAG4D (Zeng et al., 2024), 4Diffusion (Zhang et al., 2024), DreamGaussian4D (DG4D) (Ren et al., 2023), GaussianFlow (Gao et al., 2024), 4DGen (Yin et al., 2023), Efficient4D (Pan et al., 2024b). More baseline details are in Appendix C.
@@
SV3D (Voleti et al., 2024) 0.129 0.925 989.53
4Diffusion (Zhang et al., 2024) 0.164 0.863 -
Diffusion2 (Yang et al., 2025) 0.189 0.907 1205.16
STAG4D (Zeng et al., 2024) 0.131 0.929 861.88
@@
Consistent4D (Jiang et al., 2024c) 0.160 0.87 1133.93
STAG4D (Zeng et al., 2024) 0.126 0.91 992.21
4Diffusion (Zhang et al., 2024) 0.165 0.88 -
Efficient4D (Pan et al., 2024b) 0.130 0.92 -
4DGen (Yin et al., 2023) 0.140 0.89 -
DG4D (Ren et al., 2023) 0.160 0.87 -
GaussianFlow (Gao et al., 2024) 0.140 0.91 -
@@
Table 3: Evaluation of novel view video synthesis on the ObjaverseDy dataset. SV4D can achieve superior performance in both video frame and multi-view consistency. Our reproduced version.
@@
SV3D (Voleti et al., 2024) 0.131 0.920 729.67 375.49 526.78 690.49
Diffusion2 (Yang et al., 2025) 0.188 0.869 1048.47 564.80 938.98 1320.29
STAG4D (Zeng et al., 2024) 0.133 0.917 652.43 469.07 636.83 546.56
@@
Consistent4D (Jiang et al., 2024c) 0.165 0.896 880.54 488.38 741.52 871.95
STAG4D (Zeng et al., 2024) 0.158 0.860 929.10 453.62 663.50 1003.16
DreamGaussian4D (Ren et al., 2023) 0.152 0.897 697.80 450.58 615.68 638.15
@@ 4.1 Quantitative Comparison
Figure 6: Visual comparison of generated 4D outputs. We show two frames in the input videos and render two novel views of the corresponding frames. By leveraging the consistent multi-view videos generated by SV4D, we can learn higher-quality 4D assets compared to the prior works with SDS-based losses. Our results are more detailed, consistent, and faithful to the input videos.
@@ 4.2 Visual Comparison
4D Generation. We compare our generated 4D results with the prior methods in Fig. 6. For each video, we render the 4D outputs at two different timestamps and two novel views. Consistent4D and STAG4D often produce blurry outputs with inconsistent geometry and texture. DreamGaussian4D can generate finer texture details but still suffers from flickering artifacts and sometimes creates inaccurate geometry. Moreover, all these methods rely on the computationally expensive SDS loss, which is prone to spatially incoherent results and over-saturated texture. On the contrary, SV4D optimizes 4D assets using purely photometric and geometric losses, resulting in smoother videos with realistic and faithful object appearance. More visual comparison results can be found in Appendix B.1.
@@ 4.3 User Study
SV3D (Voleti et al., 2024) 700.72 831.18 656.59 780.99
Diffusion2 (Yang et al., 2025) 896.98 890.70 801.99 1093.24
STAG4D (Zeng et al., 2024) 765.79 669.28 601.23 659.79
@@ 4.4 Ablative Analyses
SV4D Sampling is better than off-the-shelf interpolation method. Extending anchor frames to the full $V{\mkern-2.0mu\times\mkern-2.0mu}F$ image matrix is not trivial. In Table 6, we compare our SV4D sampling (ours) and the off-the-shelf interpolation method AMT (Li et al., 2023). The model with SV4D sampling outperforms AMT across all metrics. Fig. 8 validates this observation, where we can see that the SV4D sampling can synthesize distinct images while the results of the off-the-shelf interpolation method have a lot of blur results or missing parts. Table 6 also shows the results of two intuitive sampling strategies, independent and interleaved sampling, both of which produce videos with lower consistency. Appendix A.6 explains the details of independent and interleaved sampling. These results validate that the SV4D sampling is effective.
@@ 5 Conclusion
Appendix A Implementation Details
A.1 Data Details
Similar to SV3D (Voleti et al., 2024), we render views of a subset of the Objaverse dataset (Deitke et al., 2023b) curated CC-licensed animatable 3D objects from the original dataset. We held out some objects for evaluation purposes.
Each loaded object is scaled such the largest world-space XYZ extent of its bounding box is 1. We use Blender’s CYCLES renderer to render multiple views and video frames for each object. We limit the number of samples in CYCLES renderer to save the rendering time. For lightning, we follow SV3D to randomly select from a set of curated HDRI envmaps.
For static orbits, we regularly sample azimuth angles with a constant value starting from azimuth 0. We randomly sample one elevation angle for all views. For dynamic orbits, the sequence of camera elevations for each orbit is obtained from a random weighted combination of sinusoids with different frequencies (Voleti et al., 2024). The azimuth angles are sampled regularly, and then a small amount of noise is added to make them irregular. The elevation values are smoothed using a simple convolution kernel and then clamped to a maximum elevation.
We then encode all of these images into latent space using SD2.1 (Rombach et al., 2022)’s VAE, and CLIP (Radford et al., 2021). We then store the latent and CLIP embeddings for all of these images along with the corresponding elevation and azimuth values, and frame index.
A.2 SV4D Training Details
Our approach involves utilizing the widely used EMD (Karras et al., 2022), incorporating a $L2$ loss for fine-tuning, as followed in SVD (Blattmann et al., 2023b) and SV3D (Voleti et al., 2024). We adopt $L2$ loss during the training. To optimize the training efficiency and reduce GPU VRAM, we follow SV3D to reprocess and store the precomputed latent and CLIP embeddings for all images in advance. During training, these tensors are directly loaded rather than being computed in real-time. We found that with 8 views and 5 frames we were able to fit a batch size of 1 at $576\times 576$ resolution. We randomly sample 8 views and 5 frames from our 21 rendered views and 21 frames for training.
A.3 SV4D Inference Details
During inference, we use 50 steps of the deterministic DDIM sampler (Song et al., 2021). We adopt the SV4D sampling to generate the full V×$\times$×F image matrix during the inference. We take $V=8$ and $F=21$ as an example. Our SV4D can generate 8 views $\times$ 5 frames in an inference. In the interleaved sampling, we first generate a sparse set of anchor frames with the frame indexes [4, 8, 12, 16, 20]. Then in the dense sampling, we use the anchor frames as new reference images to densely sample the remaining frames. For example, in the first run, we generate frames with the indexes [0, 1, 2, 3, 4] with frame 0 (forward) or frame 4 (backward) as conditioning. It takes about 40 seconds to generate 5 frames and 8 views for the SV4D model.
A.4 SV4D Network Architecture
SV4D network consists of three attention layers: spatial attention, view attention, and frame attention. We reshape the feature $\mathbf{F}$ before feeding it to each attention layer $\gamma_{(\cdot)}(\cdot)$.
The spatial attention models spatial coherence across image spatial locations by calculating the attention of points at the same image:
$\mathbf{F}=\text{rearrange}(\mathbf{F},~{}B~{}V~{}F~{}H~{}W~{}D\rightarrow(B~{}V~{}F)~{}(H~{}W)~{}D)$ (1)
$\mathbf{F}=\gamma_{s}(\mathbf{F})$ (2)
$\mathbf{F}=\text{rearrange}(\mathbf{F},~{}(B~{}V~{}F)~{}(H~{}W)~{}D\rightarrow~{}B~{}V~{}F~{}H~{}W~{}D)$ (3)
The view attention models coherence across views by calculating the attention of points at the same spatial location and the same frame across views:
$\mathbf{F}=\text{rearrange}(\mathbf{F},~{}B~{}V~{}F~{}H~{}W~{}D\rightarrow(B~{}F~{}H~{}W)~{}V~{}D)$ (4)
$\mathbf{F}=\gamma_{v}(\mathbf{F})$ (5)
$\mathbf{F}=\text{rearrange}(\mathbf{F},~{}(B~{}F~{}H~{}W)~{}V~{}D\rightarrow~{}B~{}V~{}F~{}H~{}W~{}D)$ (6)
The frame attention models coherence across frames by calculating the attention of points at the same spatial location and the same view across frames:
$\mathbf{F}=\text{rearrange}(\mathbf{F},~{}B~{}V~{}F~{}H~{}W~{}D\rightarrow(B~{}V~{}H~{}W)~{}F~{}D)$ (7)
$\mathbf{F}=\gamma_{f}(\mathbf{F})$ (8)
$\mathbf{F}=\text{rearrange}(\mathbf{F},~{}(B~{}V~{}H~{}W)~{}F~{}D\rightarrow~{}B~{}V~{}F~{}H~{}W~{}D)$ (9)
$B$ is batch size, $V$ is number of views, $F$ is number of frames. $H$, $W$, and $D$ are width, height, and feature dimension of the image feature. In our experiments, both $V$ and $F$ are 8 and 5, respectively.
Note that our frame attention layers are skip-connected to the outputs of view attention with a learnable blending weight per layer, allowing it to effectively merge the spatial and temporal information in the final output. Although an optimal way is to perform spatial-temporal attention jointly without considering memory limitation, we find that our sequential design can best leverage the priors in SVD and SV3D with minimal computation overhead.
Figure 9: Independent sampling. The submatrix sampled at each inference consists of consecutive frames and each submatrix is sampled separately, without relying on previous fully sampled images.
Figure 10: Interleaved Sampling. The submatrix sampled at each inference consists of interleaved frames and each submatrix is sampled separately, without relying on previous fully sampled images.
A.5 Dynamic NeRF Optimization Details
Our main 4D reconstruction losses are the pixel-level mean squared error $\mathcal{L}_{\text{mse}}=\lVert\bm{M}-\bm{\hat{M}}\rVert^{2}$, LPIPS (Zhang et al., 2018a) loss $\mathcal{L}_{\text{lpips}}$, and mask loss $\mathcal{L}_{\text{mask}}=\lVert\bm{S}-\bm{\hat{S}}\rVert^{2}$, where $\bm{S}$, $\bm{\hat{S}}$ are the predicted and ground-truth masks. We further employ a normal loss using the estimated mono normal by Omnidata (Eftekhar et al., 2021), which is defined as the cosine similarity between the rendered normal $\bm{n}$ and estimated pseudo ground truths $\bm{\bar{n}}$: $\mathcal{L}_{\text{normal}}=1-(\bm{n}\cdot\bm{\bar{n}})$. To regularize the output geometry, we apply a smooth depth loss inspired by RegNeRF (Niemeyer et al., 2022): $\mathcal{L}_{\text{depth}}(i,j)=\left(d(i,j)-d(i,j+1)\right)^{2}+\left(d(i,j)-(d(i+1,j)\right)^{2}$, where $i,j$ indicate the pixel coordinate. For surface normal we instead rely on a bilateral smoothness loss similar to Boss et al. (2022). We found that this is crucial to getting high-frequency details and avoiding over-smoothed surfaces. For this loss we compute the image gradients of the input image $\nabla\bm{I}$ with a Sobel filter (Kanopoulos et al., 1988). We then encourage the gradients of rendered normal $\nabla\bm{n}$ to be smooth if (and only if) the input image gradients $\nabla\bm{I}$ are smooth. The loss can be written as $\mathcal{L}_{\text{bilateral}}=e^{-3\nabla\bm{I}}\sqrt{1+||\nabla\bm{n}||}$. The overall objective is then defined as the weighted sum of these losses. All losses are applied in both static and dynamic stages. We use an Adam optimizer (Kingma & Ba, 2014) with a learning rate of $0.01$ for both stages.
A.6 Different Sampling Strategies
Due to memory limitations, we cannot generate all the novel view frames at once. To generate the full image matrix, one can run the submatrix generation in different strategies. To validate the effectiveness of our SV4D sampling, we compare ours with two other sampling strategies: independent sampling and interleaved sampling, which are shown in Fig. 9 and Fig. 10, respectively. For illustrative purposes only, the SV4D model can generate 2 frames and 2 views at each inference. In the independent sampling, the submatrix sampled at each inference consists of consecutive frames and each submatrix images are sampled separately, without relying on previous fully sampled images. In the interleaved sampling, the submatrix sampled at each inference consists of interleaved frames.
Appendix B Additional Results
Figure 11: Visual Comparison of Novel View Video Synthesis Results. We show two frames in the input videos and one novel-view result of the corresponding frames. Compared to the baseline methods, SV4D outputs contain geometry and texture details that are more faithful to the input video and consistent across frames.
Figure 12: Visual Comparison of Generated 4D Outputs. We show two frames in the input videos and render one view result of the corresponding frames. By leveraging the consistent multi-view videos generated by SV4D, we can learn higher-quality 4D assets compared to the prior works with SDS-based losses. Our results are more detailed, consistent, and faithful to the input videos.
B.1 More Visual Comparisons
We show more visual comparisons of novel view synthesis and 4D generation, as shown in Fig. 11 and Fig. 12. These results further demonstrate that SV4D-generated images and 4D results are more consistent and detailed, faithful to the conditioning videos compared to the prior works.
B.2 More Ablative Results
Table 7: Additional Ablations. SV4D w/o frame atten.: SV4D without frame attention. SV4D w/o MV: SV4D without reference multi-views. SV4D with CLIP MV: SV4D with CLIP embeddings (instead of VAE latents) of the first frame for frame attention conditioning and without reference multi-views.
[…]
SV4D w/o frame atten. 0.131 0.920 729.67 375.49 526.78 690.49
SV4D w/o MV 0.135 0.870 876.07 410.90 608.93 710.66
SV4D with CLIP MV 0.142 0.868 1174.47 530.74 819.22 1327.75
SV4D 0.136 0.920 585.09 331.94 503.02 470.46
We show an additional ablation study on the SV4D network architecture to justify our model design in Table 7. In particular, we report the quantitative comparison of ablated models: SV4D without frame attention, SV4D without multi-view conditioning, and SV4D with CLIP embeddings (instead of VAE latents) of the first frame for frame attention conditioning while removing the reference multi-views. The results justify that using VAE latents of reference multi-views as frame attention conditioning achieves better performance in most metrics compared to other design choices.
Figure 13: FVD-F vs. View Number on the ObjaverseDy dataset. We observe that SV4D produces better temporal consistency in almost all views.
B.3 Quality Per View
We plot the FVD-F value for each view on the ObjaverseDy dataset, as shown in Fig. 13. We observe that SV4D produces better temporal consistency in almost all views.
Appendix C Baseline Details
For novel view video synthesis (Table 23, and 6), we compare SV4D with SV3D (Voleti et al., 2024), Diffusion2 (Yang et al., 2025), STAG4D (Yang et al., 2025), 4Diffusion (Zhang et al., 2024). The results of SV3D, Diffusion2, and 4Diffusion are generated with their official code. STAG4D used Zero123++ (Shi et al., 2023) as the multi-view generator, which fixed the view angle, and hence the novel views generated from STAG4D cannot be changed to be consistent with the views evaluated. We reproduced STAG4D with SV3D as the multi-view generator. SV3D has been shown to generate more consistent 3D results than Zero123++, so this serves as a stronger baseline.
For 4D generation, we compare SV4D with Consistent4D (Jiang et al., 2024c), STAG4D (Zeng et al., 2024), 4Diffusion (Zhang et al., 2024), DreamGaussian4D (DG4D) (Ren et al., 2023), GaussianFlow (Gao et al., 2024), 4DGen (Yin et al., 2023), Efficient4D (Pan et al., 2024b). On the Consistent4D dataset (Table 2), the results of Consistent4D, STAG4D, Efficient4D, and GaussianFlow are from the original paper. 4DGen results are from Efficient4D paper, and DG4D results are from GaussianFlow paper. The results of 4Diffusion are generated with its official code. On the ObjaverseDy dataset (Table 4), all results of our baselines are generated from their official code.
Appendix D Supplementary Video
Beyond the paper, our supplementary materials offer a comprehensive video that provides an in-depth introduction to our task and method. Additionally, we add more visual comparisons which further showcase the effectiveness of our approach.